OpenStudy (anonymous):

If the bases are 1/2 and 1/4, and the height is 1, what is the area?

OpenStudy (anonymous):

What shape is this?

OpenStudy (anonymous):



OpenStudy (anonymous):

A = 1/2(b1 + b2)h A = 1/2(1/2 + 1/4)1 A = 1/2(3/4) A = 3/8

OpenStudy (binary3i):

 imagine two triangles.